2x-y+8z=13
3x+4y+5z=18
5x-2y+7z=20

Answer by ankor@dixie-net.com(22740)

2x-y+8z=13
3x+4y+5z=18
5x-2y+7z=20
:
Using the first two equations, multiply the 1st equation by 4, add to the 2nd
8x - 4y + 32z = 52
3x + 4y + 5z = 18
-------------------addition eliminates y
11x + 37z = 70
:
using the second and third equations, multiply the third by 2
3x + 4y + 5z = 18
10x -4y + 14z = 40
-------------------addition eliminates y
13x + 19z = 58
:
We have two 2 unknown equations
11x + 37z = 70
13x + 19z = 58
multiply the first equation by 13, multiply the second equation by 11
143x + 481z = 910
143x + 209z = 638
------------------Subtraction eliminates x find z
0 + 272z = 272
z = 1
Find x using 13x + 19z = 58
13x + 19 = 58
13x = 58 - 19
13x = 39
x = 39/3
x = 3
find y using the very first equation 2x-y+8z=13
2(3) - y + 8(1) = 13
6 - y + 8 = 13
- y = 13 - 14
-y = -1
y = 1
;
:
You can confirm this for yourself in the other two original equations